OpenStudy (anonymous):

(6 + sq rt of 7)(4 + sq rt of 7) a. 7 + 10 sq rt 7 b. 31 c. 31+10 sq rt 7 d. 73

Parth (parthkohli):

The best and the only thing you do here is FOIL it.

OpenStudy (lgbasallote):

hmm...so where are you confused here??

Parth (parthkohli):

\(\Rightarrow 6(4 + \sqrt7) + \sqrt7(4 + \sqrt7) \)

Parth (parthkohli):

Do you know distribution?

OpenStudy (anonymous):

A

OpenStudy (lgbasallote):

i thought you said FOIL lol =))

OpenStudy (anonymous):

\[(a+b)(c+d) = a(c+d) + b(c+d) = ac + ad + bc + bd\] a = 6 b = sqrt(7) c = 4 d = sqrt(7)

Parth (parthkohli):

Well, can you work it out for us?

OpenStudy (anonymous):

The answer is not A. Please show your work so we can help you understand. If you use the equation I gave, it should walk you right through it.

Parth (parthkohli):

I got a bug! It says that I got a medal but I didn't

OpenStudy (anonymous):

Not a bug. They took it back.

Parth (parthkohli):

I see.

OpenStudy (anonymous):

i got it guys (: nvm.